# Graph-Structured Topic Modeling for Documents with Spatial or Covariate Dependencies

### v1

@@ B.2 Analysis of the Estimation of $W$
In this section, we adapt the proof of Klopp et al. (2021) that derives a high probability bound for the outcome $\widehat{W}$ after successive projections. We evaluate the vertices $\widehat{H}$ detected by SPA with the rows of $\widehat{U}$ as the input. To accomplish this, we first need to bound the row-wise error of $\widehat{U}$ which is closely related to the upper bound of the estimated vertices $\widehat{H}=\widehat{U}_{J}$ and ultimately, is linked to $\widehat{W}=\widehat{U}\widehat{H}^{-1}$. Similarly to Klopp et al. (2021), define $\beta(M,X)$ as,
$\max_{i=1,\cdots,n}\|\mathbf{e}_{i}^{T}(\widehat{U}-UO)\|_{2}\leq\beta(M,X)$ (40)
We first need the following assumption on $\beta(M,X)$.
Assumption 6
@@
$\beta(M,X)\leq\frac{\bar{C}}{\lambda_{1}(W)\kappa(W)K\sqrt{K}}$
Lemma 2 (Adapted from Corollary 5 of Klopp et al. (2021))
Let Assumptions 1 to 6 hold. Assume that $M\in\mathbb{R}^{n\times p}$ is a rank-$K$ matrix. Let $U,\widehat{U}\in\mathbb{R}^{n\times K}$ be the left singular vectors corresponding to the top $K$ singular values of $M$ and its perturbed matrix $X\in\mathbb{R}^{n\times p}$, respectively. Let $J$ be the set of indices returned by the SPA with input $(\widehat{U},K)$, and $\widehat{H}=\widehat{U}_{J}$. Let $O\in\mathbb{O}_{K}$ be the same matrix as (13). Then, for a small enough $\bar{C}$, there exists a constant $C^{{}^{\prime}}>0$ and a permutation $\tilde{P}$ such that
$\|\widehat{H}-\tilde{P}HO\|_{F}\leq C^{{}^{\prime}}\sqrt{K}\kappa(W)\beta(M,X)$ (41)
where $\beta(M,X)$ is defined as (40).
Proof  The proof here is a direct combination of Theorem 4 and Corollary 5 of Klopp et al. (2021), for SPA (rather than pre-conditioned SPA). The crux of the argument consists of applying Theorem 2 in (Gillis and Vavasis, 2015) which bounds the error of SPA in the near-separable nonnegative matrix factorization setting,
@@
where $Q=HO$ and $N\in\mathbb{R}^{n\times K}$ is the noise matrix. Note that the rows of $U$ lie on a simplex with vertices $H$ and weights $W$. The theorem states that if the row-wise error of $N$ is bounded by an $\epsilon$ such that:
$\|\mathbf{e}_{i}^{\top}N\|_{2}\leq\epsilon\leq C_{*}\frac{\lambda_{\min}(Q)}{K\sqrt{K}}$ (42)
for some small constant $C_{*}>0$, we can bound the column-wise error of $\widehat{U}$ for columns in the set $J$ under permutation $\pi$,
$\|\widehat{U}_{j}-q_{\pi(j)}\|_{2}\leq C^{{}^{\prime}}\kappa(Q)\epsilon$
Applying this result to our setting, under Assumption 4,6 and Lemma 5, our bound on $\|\mathbf{e}_{i}^{\top}N\|_{2}=\|\mathbf{e}_{i}^{\top}(\widehat{U}-UO)\|_{2}$ satisfies (42),
@@
for a small enough $\bar{C}\leq C_{*}$. Noting that $\widehat{H}=\widehat{U}_{J}$ and $\kappa(H)=\kappa(W)$, with the permutation matrix $\tilde{P}$ corresponding to $\pi$, we get
$\|\widehat{H}-\tilde{P}HO\|_{F}\leq C^{{}^{\prime}}\sqrt{K}\kappa(W)\beta(M,X)$ (43)
@@
Lemma 3 (Adapted from Lemma 2 of Klopp et al. (2021))
@@
$\min_{P\in\mathcal{P}}\|\widehat{W}-WP\|_{F}\leq 2C^{{}^{\prime}}\sqrt{K}\lambda_{1}^{2}(W)\kappa(W)\beta(X,\Gamma)+\lambda_{1}(W)\|\widehat{U}-UO\|_{F}$
@@
Proof  The first part of the proof on the invertibility of $\widehat{H}$ is analogous to Lemma 2 in Klopp et al. (2021), where combined with Lemma 5, we obtain the inequality,
@@
$\begin{split}\|\widehat{W}-WP\|_{F}&=\|\widehat{U}\widehat{H}^{-1}-UH^{-1}P\|_{F}\\ &\leq\|\widehat{U}(\widehat{H}^{-1}-O^{\top}H^{-1}P)\|_{F}+\|(\widehat{U}-UO)[P^{-1}HO]^{-1}\|_{F}\\ &\leq\|\widehat{H}^{-1}\|_{op}\|H^{-1}\|_{op}\|\widehat{H}-\tilde{P}HO\|_{F}+\|\widehat{U}-UO\|_{F}\|H^{-1}\|_{op}\\ &\leq 2C^{{}^{\prime}}\sqrt{K}\lambda_{1}^{2}(W)\kappa(W)\beta(X,\Gamma)+\lambda_{1}(W)\|\widehat{U}-UO\|_{F}\end{split}$
where we used the well known inequality $\|A^{-1}-B^{-1}\|_{F}\leq\|A^{-1}\|_{op}\|B^{-1}\|_{op}\|A-B\|_{F}$.
B.2.1 Proof of Theorem 3
@@
Proof  Notably, by the property of maximum row norms, $\max_{i=1,\cdots,n}\|\mathbf{e}_{i}^{\top}(\widehat{U}-UO)\|_{2}\geq\|\widehat{U}-UO\|_{F}/\sqrt{n}$. It is then sufficient to set $\beta(M,X)$ defined in (40),
$\beta(M,X)=\frac{10C}{\sqrt{n}\lambda_{K}(M)}\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)$ (44)
which is a high probability bound on $\max_{i=1,\cdots,n}\|\mathbf{e}_{i}^{\top}(\widehat{U}-UO)\|_{2}$ with probability at least $1-o(n^{-1})$ (Theorem 2).
@@
$\lambda_{k}(M)\geq c\lambda_{1}(W)\geq c\sqrt{n/K}$ (45)
Then using the condition on $N$, (12),
@@
$\begin{split}\|\sin\Theta(V,\widehat{V}^{0})\|_{F}&\leq\frac{C}{\lambda_{K}(M)^{2}}K\sqrt{\frac{n\log(n)}{N}}\\ &\leq\frac{CK^{2}}{c^{2}n}\sqrt{\frac{n\log(n)}{N}}\\ &\leq\frac{C}{c^{2}(\sqrt{n_{\mathcal{C}}+\rho^{2}(\Gamma)s\lambda_{\max}(\Gamma)})}\leq\frac{1}{2}\end{split}$ (46)
for large enough $c>0$.
Next from (12), (44), and (45),
$\beta(M,X)\leq\frac{1}{\sqrt{c^{*}_{\min}}K\sqrt{K}\lambda_{K}(M)}\leq\frac{\bar{C}}{\lambda_{1}(W)\kappa(W)K\sqrt{K}}$ (47)
@@
which proves Assumption 6. Thus, we are ready to use Theorem 2 and Lemma 3. We can now plug in $\beta(M,X)$ (44) and the error bound of graph-regularized SVD (13) in the bound of $\min_{P\in\mathcal{P}}\|\widehat{W}-WP\|_{F}$ in Lemma 3.
@@
$\begin{split}\|\widehat{W}-WP\|_{F}&\leq 2C^{{}^{\prime}}\sqrt{K}\lambda_{1}^{2}(W)\kappa(W)\beta(X,\Gamma)+\lambda_{1}(W)\|\widehat{U}-UO\|_{F}\\ &\leq\frac{20C^{{}^{\prime}}C}{\sqrt{n}}\left(\frac{\lambda_{1}(W)}{\lambda_{K}(M)}\right)^{2}\lambda_{K}(M)\kappa(W)K\sqrt{\frac{\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\\ &\qquad+10C\frac{\lambda_{1}(W)}{\lambda_{K}(M)}\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\\ &\leq\frac{20c^{*}C^{{}^{\prime}}C}{\sqrt{n}c^{2}}\lambda_{K}(M)K\sqrt{\frac{\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\\ &\qquad+\frac{10C}{c}\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\\ &\leq\frac{20C}{c}K\sqrt{\frac{\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\end{split}$ (48)
Here, we used the bounds on condition numbers in Assumption 4 and $\lambda_{K}(M)\leq\sqrt{n}$ in Lemma 4.
B.3 Analysis of the Estimation of $A$
Using the result of Theorem 3, we now proceed to bound the error of $\widehat{A}$ in Step 4 (9).
Proof  By the simple basic inequality, letting $P=\arg\min_{O\in\mathcal{F}}\|\widehat{W}-WO\|_{F}$:

### v2

@@ Appendix D Analysis of the Estimation of $W$ and $A$
In this section, we adapt the proof of Klopp et al. (2021) that derives a high probability bound for the outcome $\widehat{W}$ after successive projections. We evaluate the vertices $\widehat{H}$ detected by SPA with the rows of $\widehat{U}$ as the input. To accomplish this, we first need to bound the row-wise error of $\widehat{U}$ which is closely related to the upper bound of the estimated vertices $\widehat{H}=\widehat{U}_{J}$ and ultimately, is linked to $\widehat{W}=\widehat{U}\widehat{H}^{-1}$.
To apply Theorem 1 of Gillis & Vavasis (2015) on the estimation with SPA, we need to show that the error on each of the row of the estimated left singular vector of $MV^{\top}=U\Lambda$ is controlled, which requires us bounding the error: $\|\widehat{U}-UO\|_{2\to\infty}=\max_{i\in[n]}\|e_{i}^{\top}(\widehat{U}-UO)\|_{2}$.
Lemma 2 (Baseline two-to-infinity norm bound (Theorem 3.7 of Cape et al. (2019)).
For $C,E\in\mathbb{R}^{n\times p}$, denote $\widehat{C}:=C+E$ as the observed matrix that adds perturbation $E$ to unobserved $C$. For $C$ and $\widehat{C}$, their respective singular value decompositions of are given as
$\begin{split}C&=U\Lambda V^{\top}+U_{\perp}\Lambda_{\perp}V_{\perp}^{\top}\\ \widehat{C}&=\widehat{U}\widehat{\Lambda}\widehat{V}^{\top}+\widehat{U}_{\perp}\widehat{\Lambda}_{\perp}\widehat{V}_{\perp}^{\top}\end{split}$
where $\Lambda,\widehat{\Lambda}\in\mathbb{R}^{r\times r}$ contain the top $r$ singular values of $C,\widehat{C}$, while $\Lambda_{\perp},\widehat{\Lambda}_{\perp}\in\mathbb{R}^{n-r\times p-r}$ contain the remaining singular values. Provided $\lambda_{r}(C)>\lambda_{r+1}(C)\geq 0$ and $\lambda_{r}(C)\geq 2\|E\|_{op}$, then,
$\|\hat{U}-UW_{U}\|_{2\to\infty}$ $\leq 2\left(\frac{\|(U_{\perp}U_{\perp}^{\top})E(VV^{\top})\|_{2\to\infty}}{\sigma_{r}(C)}\right)$
$\quad+2\left(\frac{\|(U_{\perp}U_{\perp}^{\top})E(V_{\perp}V^{\top})\|_{2\to\infty}}{\sigma_{r}(C)}\right)\|\sin\Theta(\hat{V},V)\|_{op}$
$\quad+2\left(\frac{\|(U_{\perp}U_{\perp}^{\top})C(V_{\perp}V^{\top})\|_{2\to\infty}}{\sigma_{r}(C)}\right)\|\sin\Theta(\hat{V},V)\|_{op}$
$\quad+\|\sin\Theta(\hat{U},U)\|_{op}^{2}\|U\|_{2\to\infty}.$ (43)
where $W_{U}$ is the solution of $\inf_{W\in\mathbb{O}_{r}}\|\widehat{U}-UW\|_{F}$.
Proof.
The proof is given in Section 6 of Cape et al. (2019). ∎
Adapting Lemma 2 to our setting, we set $C=U\Lambda$ and $\widehat{C}=\bar{U}^{t_{\max}}$. We also use the notation $C^{*}$ to denote the oracle, $C^{*}=U\Lambda V^{T}\hat{V}_{t_{\max}-1}.$ We set $r=K$ and denote the $SVD$s of $C^{*}$ and $\hat{C}$:
Note that we are performing a rank $K$ decomposition, so in the previous SVD, $V_{\perp}=0$, by which (2) becomes,
$\begin{split}\|\hat{U}-UO\|_{2\to\infty}&\leq 2\left(\frac{\|(U_{\perp}U_{\perp}^{\top})E(VV^{\top})\|_{2\to\infty}}{\lambda_{K}(M)}\right)\\ &\quad+\|\sin\Theta(\hat{U},U)\|_{op}^{2}\|U\|_{2\to\infty}.\end{split}$
with $E=\hat{C}-C^{*}+C^{*}-C.$ We thus need to bound the quantity,
$\|(U_{\perp}U_{\perp}^{\top})E(VV^{\top})\|_{2\to\infty}\leq\|P_{U^{\perp}}E\|_{2,\infty}\leq\underbrace{\|P_{U^{\perp}}(\hat{C}-C^{*})\|_{2,\infty}}_{(A)}+\underbrace{\|P_{U^{\perp}}(C^{*}-C)\|_{2,\infty}}_{(B)}.$
The second term (B) is 0 because $P_{U^{\perp}}(C^{*}-C)=P_{U^{\perp}}(U\Lambda V^{T}\widehat{V}_{t_{\max}-1}-U\Lambda)=P_{U^{\perp}}U\Lambda(V^{T}\widehat{V}_{t_{\max}-1}-I_{K})=0$. For the first term (A), we note that $\hat{U}$ stems from the SVD of $\widehat{C}=\bar{U}_{t_{\max}}$, which is itself the solution of:
$\widehat{C}=\text{argmin}_{C\in\mathbb{R}^{n\times K}}\frac{1}{2}\|C-\tilde{Y}\|_{F}^{2}+\rho\sum_{i\sim j}\|C_{i\cdot}-C_{j\cdot}\|_{2}$
where $\tilde{Y}=X\widehat{V}_{t_{\max}-1}$. By the KKT conditions, for any $i\in[n]$,
$\widehat{C}_{i\cdot}-\tilde{Y}_{i\cdot}+\rho\sum_{j\sim i}z_{ij}=0$
where $z_{ij}$ denotes the subgradient of $\|C_{i\cdot}-C_{j\cdot}\|_{2}$ so that $z_{ij}(k)=\frac{\hat{C}_{ik}-\hat{C}_{jk}}{\|\hat{C}_{i}-\hat{C}_{j}\|_{2}}$ and $\|z_{ij}\|_{2}<1$ if $\hat{C}_{i\cdot}=\hat{C}_{j\cdot}$ and $\|z_{ij}\|_{2}=1$ if $\hat{C}_{i}-\hat{C}_{j}\neq 0.$
Therefore:
$\begin{split}\hat{C}_{i\cdot}-C^{*}_{i\cdot}&=\tilde{Y}_{i\cdot}-C^{*}_{i\cdot}+\rho\sum_{j\sim i}z_{ij}\\ \implies\|\hat{C}_{i\cdot}-C^{*}_{i\cdot}\|_{2}&\leq\|\tilde{Y}_{i\cdot}-C_{i\cdot}^{*}\|_{2}+\rho\sum_{j\sim i}\|z_{ij}\|_{2}\\ &\leq\|\tilde{Y}_{i\cdot}-C_{i\cdot}^{*}\|_{2}+\rho d_{\max}\\ \implies\|\hat{C}-C^{*}\|_{2\to\infty}&\leq\|Z\hat{V}_{t_{\max}-1}\|_{2\to\infty}+\rho d_{\max}.\end{split}$ (44)
We have:
$\begin{split}\|Z\hat{V}_{t_{\max}-1}\|_{2\to\infty}&=\|Z(VV^{T}+V_{\perp}V_{\perp}^{T})\hat{V}_{t_{\max}-1}\|_{2\to\infty}\\ &\leq\|ZVV^{T}\hat{V}_{t_{\max}-1}\|_{2\to\infty}+\|ZV_{\perp}V_{\perp}^{T}\hat{V}_{t_{\max}-1}\|_{2\to\infty}\\ &\leq\|ZV\|_{2\to\infty}\|V^{T}\hat{V}_{t_{\max}-1}\|_{op}+\|ZV_{\perp}\|_{2\to\infty}\|V_{\perp}^{T}\hat{V}_{t_{\max}-1}\|_{op}\\ \end{split}$ (45)
Then, by Lemma 17, we have
$\begin{split}\|Z\hat{V}_{t_{\max}-1}\|_{2\to\infty}&\leq\|ZV\|_{2\to\infty}+\|Z\|_{2\to\infty}L_{t_{\max-1}}\\ &\leq C_{1}\sqrt{\frac{K\log(n)}{N}}+C_{2}\sqrt{\frac{K\log(n)}{N}}L_{t_{\max}-1}\\ \end{split}$
Using the derivation of the geometric series of errors in (36), recall
$\begin{split}u&=\frac{5/2C}{\lambda_{K}(M)}\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\\ r&=\frac{200/49C}{\lambda_{K}(M)}\sqrt{\frac{K\log(n)}{N}}\left(\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\vee\sqrt{n}\right)\end{split}$
and expressing $L_{t_{\max}-1}$ in terms of $u$ and $v$,
$\begin{split}\|Z\hat{V}_{t_{\max}-1}\|_{2\to\infty}&\leq C_{1}\sqrt{\frac{K\log(n)}{N}}+C_{2}\sqrt{\frac{K\log(n)}{N}}\cdot u\frac{1-r^{t_{\max}-1}}{1-r}+\sqrt{\frac{K\log(n)}{N}}\cdot r^{t_{\max}-1}L_{0}\\ &\leq C_{1}\sqrt{\frac{K\log(n)}{N}}+C_{2}\sqrt{\frac{K\log(n)}{N}}\frac{u}{1-r}+\sqrt{K}r^{t_{\max}}L_{0}\\ &\leq C_{3}\sqrt{\frac{K\log(n)}{N}}\frac{u}{1-r}\qquad\text{since }\frac{u}{1-r}\geq r^{t_{\max}}L_{0}\\ &\leq\frac{C^{{}^{\prime}}}{\lambda_{K}(M)}\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\end{split}$
Plugging this into (D),
$\begin{split}\|\hat{U}-UO\|_{2\to\infty}&\leq 2\left(\frac{\|Z\hat{V}_{t_{\max}-1}\|_{2\to\infty}+\rho d_{\max}}{\lambda_{K}(M)}\right)+\|\sin\Theta(\hat{U},U)\|_{op}^{2}\|U\|_{2\to\infty}\\ &\leq\frac{C}{\lambda_{K}^{2}(M)}\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\end{split}$
when $\rho$ is a small value that satisfies $\rho\leq\frac{1}{\lambda_{K}(M)}\sqrt{\frac{K\log(n)}{N}}\cdot\frac{\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}}{d_{\max}}$, and we know $\|\sin\Theta(\hat{U},U)\|_{op}^{2}\|U\|_{2\to\infty}\leq\|\sin\Theta(\hat{U},U)\|_{op}^{2}\leq\|\sin\Theta(\hat{U},U)\|_{op}$.
D.1 Deterministic Bounds
First, denote $\beta(M,\Gamma)$ as:
$\beta(M,\Gamma):=\frac{C}{\lambda_{K}^{2}(M)}\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)$ (46)
which is the upper bound on the maximum row error, $\max_{i=1,\cdots,n}\|{e}_{i}^{T}(\widehat{U}-UO)\|_{2}$ by (D). We need the following assumption on $\beta(M,\Gamma)$.
Assumption 6.
@@ D.1 Deterministic Bounds
$\beta(M,\Gamma)\leq\frac{\bar{C}}{\lambda_{1}(W)\kappa(W)K\sqrt{K}}$
We will show in (50) that this assumption holds in fact with high probability. Similar to Klopp et al. (2021), the proof of the consistency of our estimator relies on the following result from Gillis & Vavasis (2015).
Lemma 3 (Robustness of SPA (Theorem 1 of Gillis & Vavasis (2015)).
Let $M=WQ\in\mathbb{R}^{n\times K}$ where $Q\in\mathbb{R}^{K\times K}$ is non degenerate, and $W=[I_{r}|\tilde{W}^{\top}]^{\top}\in\mathbb{R}_{+}^{n\times K}$ is such that the sum of the entries of each row of $W$ is at most one. Let $\tilde{M}$ denote a perturbed version of $M$, with $\tilde{M}=M+N$, with:
$\|N_{j\cdot}\|_{2}=\|e_{j}^{T}N\|_{2}=\|\tilde{M}_{j\cdot}-M_{j\cdot}\|\leq\epsilon\text{ for all j}.$
Then, if $\epsilon$ is such that:
$\|{e}_{i}^{\top}N\|_{2}\leq\epsilon\leq C_{*}\frac{\lambda_{\min}(Q)}{\sqrt{K}\kappa^{2}(Q)}$ (47)
for some small constant $C_{*}>0$, then SPA identifies the rows of $Q$ up to error $O(\epsilon\kappa^{2}(Q))$, that is, the index set $J$ of vertices identified by SPA verifies:
$\max_{j\in J}\min_{\pi\in\mathcal{P}_{K}}\|\tilde{M}_{j\cdot}-Q_{\pi(j)\cdot}\|_{2}\leq C^{{}^{\prime}}\kappa^{2}(Q)\epsilon.$
The notation $\kappa(Q)=\frac{\lambda_{\max}(Q)}{\lambda_{\min}(Q)}$ is the condition number of $Q,$ and $\mathcal{P}_{K}$ denotes the set of all permutations of the set $[K]$.
Lemma 4 (Adapted from Corollary 5 of Klopp et al. (2021)).
Let Assumptions 1 to 6 hold. Assume that $M\in\mathbb{R}^{n\times p}$ is a rank-$K$ matrix. Let $U,\widehat{U}\in\mathbb{R}^{n\times K}$ be the left singular vectors corresponding to the top $K$ singular values of $M$ and its perturbed matrix $X\in\mathbb{R}^{n\times p}$, respectively. Let $J$ be the set of indices returned by the SPA with input $(\widehat{U},K)$, and $\widehat{H}=\widehat{U}_{J}$. Let $O\in\mathbb{O}_{K}$ be the same matrix as in (13) of the main manuscript. Then, for a small enough $\bar{C}$, there exists a constant $C^{{}^{\prime}}>0$ and a permutation $\tilde{P}$ such that,
$\|\widehat{H}-\tilde{P}HO\|_{F}\leq C^{{}^{\prime}}\sqrt{K}\kappa(W)\beta(M,\Gamma)$ (48)
where $\beta(M,\Gamma)$ is defined as (46).
Proof.
The proof here is a direct combination of Lemma 3 and Corollary 5 of Klopp et al. (2021), for SPA (rather than pre-conditioned SPA). The crux of the argument consists of applying Theorem 1 in (Gillis & Vavasis 2015) (rewritten in a format more amenable to our setting in Lemma 3) which bounds the error of SPA in the near-separable nonnegative matrix factorization setting,
@@
where $Q=HO$ and $N\in\mathbb{R}^{n\times K}$ is the noise matrix. Note that the rows of $U$ lie on a simplex with vertices $Q$ and weights $W$. We apply Lemma 3 with $Q=HO$, and $N=\widehat{U}-UO$.
Under Assumption 4,6 and Lemma 7, the error $\|{e}_{i}^{\top}N\|_{2}=\|{e}_{i}^{\top}(\widehat{U}-UO)\|_{2}$ satisfies:
@@
for a small enough $\bar{C}\leq C_{*}$. Thus the condition on the noise (Equation (47)) in Theorem 3 is met. Noting that $\widehat{H}=\widehat{U}_{J}$ and $\kappa(H)=\kappa(W)$, with the permutation matrix $\tilde{P}$ corresponding to $\pi$, we get
$\|\widehat{H}-\tilde{P}HO\|_{F}\leq C^{{}^{\prime}}\kappa^{2}(W)\beta(M,\Gamma)$
@@
Lemma 5 (Adapted from Lemma 2 of Klopp et al. (2021)).
@@
$\min_{P\in\mathcal{P}}\|\widehat{W}-WP\|_{F}\leq 2C^{{}^{\prime}}\sqrt{K}\lambda_{1}^{2}(W)\kappa(W)\beta(M,\Gamma)+\lambda_{1}(W)\|\widehat{U}-UO\|_{F}$
@@
Proof.
The first part of the proof on the invertibility of $\widehat{H}$ is analogous to Lemma 2 in Klopp et al. (2021), where combined with Lemma 7, we obtain the inequality,
@@
$\begin{split}\|\widehat{W}-WP\|_{F}&=\|\widehat{U}\widehat{H}^{-1}-UH^{-1}P\|_{F}\\ &\leq\|\widehat{U}(\widehat{H}^{-1}-O^{\top}H^{-1}P)\|_{F}+\|(\widehat{U}-UO)[P^{-1}HO]^{-1}\|_{F}\\ &\leq\|\widehat{H}^{-1}\|_{op}\|H^{-1}\|_{op}\|\widehat{H}-\tilde{P}HO\|_{F}+\|\widehat{U}-UO\|_{F}\|H^{-1}\|_{op}\\ &\leq 2C^{{}^{\prime}}\sqrt{K}\lambda_{1}^{2}(W)\kappa(W)\beta(M,\Gamma)+\lambda_{1}(W)\|\widehat{U}-UO\|_{F}\end{split}$
where we used the well known inequality $\|A^{-1}-B^{-1}\|_{F}\leq\|A^{-1}\|_{op}\|B^{-1}\|_{op}\|A-B\|_{F}$. ∎
Proof of Theorem 3
@@
Proof.
@@
$\lambda_{k}(M)\geq c\lambda_{1}(W)\geq c\sqrt{n/K}$
Then using the condition on $N$ (Equation 12), in Theorem 2,
@@
$\begin{split}\|\sin\Theta(V,\widehat{V}^{0})\|_{F}&\leq\frac{C}{\lambda_{K}(M)^{2}}K\sqrt{\frac{n\log(n)}{N}}\\ &\leq\frac{CK^{2}}{c^{2}n}\sqrt{\frac{n\log(n)}{N}}\\ &\leq\frac{C}{c^{2}\sqrt{c^{*}_{\min}}(\sqrt{n_{\mathcal{C}}+\rho^{2}(\Gamma)s\lambda_{\max}(\Gamma)})}\leq\frac{1}{2}\end{split}$ (49)
Next from the condition on $N$ in Theorem 3 (Equation (46)), and Assumption 4,
$\beta(M,\Gamma)\leq\frac{C\sqrt{n}}{\sqrt{c^{*}_{\min}}K\sqrt{K}\lambda_{K}^{2}(M)}\leq\frac{\bar{C}}{\lambda_{1}(W)\kappa(W)K\sqrt{K}}$ (50)
@@
which proves Assumption 6. Thus, we are ready to use Theorem 2 and Lemma 5. We can now plug in $\beta(M,\Gamma)$, the result of Equation (46) and the error bound of graph-regularized SVD (Equation (13) in Theorem 2) in the upper bound of $\min_{P\in\mathcal{P}}\|\widehat{W}-WP\|_{F}$ formulated in Lemma 5.
@@
$\begin{split}\|\widehat{W}-WP\|_{F}&\leq 2C^{{}^{\prime}}\sqrt{K}\lambda_{1}^{2}(W)\kappa(W)\beta(M,\Gamma)+\lambda_{1}(W)\|\widehat{U}-UO\|_{F}\\ &\leq\frac{2C^{{}^{\prime}}C_{1}}{\sqrt{n}}\left(\frac{\lambda_{1}(W)}{\lambda_{K}(M)}\right)^{2}\kappa(W)K\sqrt{\frac{\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\\ &\qquad+10C_{2}\frac{\lambda_{1}(W)}{\lambda_{K}(M)}\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\\ &\leq\frac{2c^{*}C^{{}^{\prime}}C_{1}}{c^{2}}K\sqrt{\frac{\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\\ &\qquad+\frac{10C_{2}}{c}\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\\ &\leq\frac{20C}{c}K\sqrt{\frac{\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\end{split}$
Here, we used the bounds on condition numbers in Assumption 4.
Proof of Theorem 4
Using the result of Theorem 3, we now proceed to bound the error of $\widehat{A}$.
Proof.
By the simple basic inequality, letting $P=\arg\min_{O\in\mathcal{F}}\|\widehat{W}-WO\|_{F}$, we get
$\begin{split}\|X-\widehat{W}\widehat{A}\|_{F}^{2}&\leq\|X-\widehat{W}P^{-1}A\|_{F}^{2}\\ \|WA+Z-\widehat{W}\widehat{A}\|_{F}^{2}&\leq\|WA+Z-\widehat{W}P^{-1}A\|_{F}^{2}\\ \|(W-\widehat{W}P^{-1})A+\widehat{W}(P^{-1}A-\widehat{A})+Z\|_{F}^{2}&\leq\|(W-\widehat{W}P^{-1})A+Z\|_{F}^{2}\end{split}$
which leads us to,